OpenStudy (anonymous):

What is the Derivative of t^(3/2) (2+t^(1/2)) Answer is 3t^(1/2)+2t I can't seem to get it.

OpenStudy (anonymous):

=3/2xt^1/2x(2+t^1/2)+t^3/2x(1/2xt^-1/2) =3t^1/2+3/2t+1/2t =3t^1/2+2t